OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 5 0 2 Line 2 x y 0 5 4 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (danjs):

the slope of each is line 1 --- m = -3/4 line 2 --- m = -3/4 the two are parallel from that

OpenStudy (danjs):

but in the table for each , when x is zero, they have different y values, 2 parallel lines, they have no solutions

OpenStudy (anonymous):

so a

OpenStudy (danjs):

y = -3/4 x + 2 y = -3/4 x + 5

OpenStudy (anonymous):

how about these Line 1 x y –4 8 4 6 Line 2 x y –1 1 3 5

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

can you put the equation for each of them together ?

OpenStudy (anonymous):

what do u mean

OpenStudy (danjs):

you get 2 points on each line, what is the equation of each line?

OpenStudy (anonymous):

I didnt understand this lesson at all

OpenStudy (danjs):

i think construction the equation for a line was before the lesson about determining how many solutions a system of 2 equations has

OpenStudy (anonymous):

not for my k12 online school

OpenStudy (danjs):

given 2 points, calculate the slope, and use y - y1 = m*(x - x1) to make the line, (x1,y1) is any point you know on the line

OpenStudy (anonymous):

im so confused but ill try

OpenStudy (anonymous):

line one the slope is - 0.25 line 2 the slope is 1

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

k...

OpenStudy (danjs):

Line 1 slope = (6-8)/(4-(-4)) = -2/8 = -1/4 Line 2 Slope= (5-1) / (3 - (-1)) = 4/4 = 1

OpenStudy (anonymous):

I think its b but im not quite sure

OpenStudy (danjs):

good, then use the slope and one of the points on the line, and make the point slope equation Line 1 y - y1 = m*( x- x1) y - 6 = (-1/4)*(x-4)

OpenStudy (danjs):

Line 2 y - 5 = 1 *(x-3)

OpenStudy (danjs):

what do we have to find again>?

OpenStudy (danjs):

oh yeah , how many solutions for the system

OpenStudy (anonymous):

so b right???

OpenStudy (danjs):

a simple thing to notice, is the two lines have different slopes they must intersect then somewhere that is a solution

OpenStudy (danjs):

what is b)

OpenStudy (anonymous):

so they would have one solution b

OpenStudy (danjs):

yeah , 2 lines with different slopes are not parallel, so they will cross eventually somewhere

OpenStudy (anonymous):

the same answers as the last question sorry

OpenStudy (danjs):

2 lines will either --intersect at some point (x,y) and have one solution to the system --Be parallel and never intersect, no solutions for this case -- be the same line , one as just a multiple of the other, infinite solutions in this case